40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4546     Entry time: Wed Apr 20 20:42:55 2011
Author: kiwamu 
Type: Update 
Category: VIDEO 
Subject: today's video session 

[Steve / Suresh / Kiwamu]

90 % of unused video cables have been removed.

Still a couple of video cables are floating around the video MUX. They will be removed in the next week's session.

DSC_2931_ss.jpg

DSC_2938_ss.jpg

 

ELOG V3.1.3-